OpenStudy (anonymous):

A positive integer is divided by a smaller positive integer. The integer part of the quotient is 9 and the remainder is 1. The sum of the two integers is 71. Find the two integers

OpenStudy (harsimran_hs4):

let the two integers be a and b (assuming a being the bigger no) so, a+b = 71 --------------------(1) and a = 9*b + 1 ---------------(2) clear till now ?

OpenStudy (anonymous):

so far so good

OpenStudy (harsimran_hs4):

great, so a = 71 - b (from (1)) not sunstitute this value of a in (2) so you get, 71 - b = 9*b + 1 now can you find b ?

OpenStudy (anonymous):

got it

OpenStudy (harsimran_hs4):

so what value of b did you get ?

OpenStudy (anonymous):

7

OpenStudy (harsimran_hs4):

good, a = 71 - b a = 71 - 7 a = 64

OpenStudy (anonymous):

Solve the following equations for a and b.\[\left\{\frac{a}{b}=9+\frac{1}{b},a+b=71\right\}\]\[\{a=64,b=7\} \]